OpenStudy (anonymous):

If tanA=2/3 and sinB =5/√41 and angles A and B are in Quadrant I, find the value of sin (A+B) and Cos (A-B). Is it possible to get a step by step?

OpenStudy (campbell_st):

Well sin(A + B) = sin(A)cos(B) + cos(A)sin(B) well to start you need 2 right triangles  and cos(A-B) = cos(A)cos(B) + sin(A)sin(B) make the necessary substitutions and evaluate